# Morphlux: Transforming Torus Fabrics for Efficient Multi-tenant ML

### v2

@@ 1. Introduction
1.1. Overcoming Semiconductor Packaging Barriers
@@ 2. Torus-based datacenter fabrics for ML
Advantages of torus fabrics. In these datacenters, accelerators are directly connected to each other in a torus-shaped topology, avoiding packet switches. The resulting direct-connect torus-based fabrics offer two key advantages. First, direct or point-to-point connectivity between accelerators eliminates delays due to network contention if the fabric topology matches the communication pattern of the collective communication primitive. And second, well-known collective communication algorithms offer optimal performance on certain contention-free torus topologies (Wu et al., 2023; Jouppi et al., 2023; Zhao et al., 2024; Sack and Gropp, 2015a). For example, the torus topology works well with multi-dimensional bucket ring algorithms (Sack and Gropp, 2015a; Jouppi et al., 2023; Jain and Sabharwal, 2010; Sack and Gropp, 2015b). This algorithm sequentially executes data transfers in a ring across all the dimensions of the torus ($XYZ$). Ring-based algorithms have an accelerator communicate with only two other accelerators at a given time, making communication in a ring on a direct-connect torus congestion-free.
@@ 3.1. L1: Bandwidth under-utilization
In a direct-connect torus topology, congestion occurs when multiple data transfers occur simultaneously on the same link (Leighton et al., 1994; Rothvoss, 2012). Accelerators in a slice can fully use their links in a dimension only when there is no congestion in that dimension. By design, an accelerator in a 3D torus has congestion-free access to links in all three dimensions only if the slice spans an entire rack or multiple racks. Consider Figure 3(a) with multiple sub-rack slices. Notice the links between slices $4$ and $3$ in the Z dimension. If both tenant slices use these links, they will cause congestion on them. To avoid causing congestion, these slices must avoid using them. As a result, the bandwidth of accelerators in each slice is underutilized by 33% because the slices have access to only 2 of the 3 dimensions. Smaller slices can have up to 66% lower bandwidth, as we show with examples in the Appendix A.
@@ 3.2. L2: Compute fragmentation
Specifically, allocations and deallocations of sub-block slices across the photonic datacenter fabric over time leads to fragmentation of resources within a rack. We quantify the fragmentation index $I$ of a rack as: $I=1-\frac{S}{T}$, where $S$ is the number of XPUs in the largest allocatable slice, and $T$ denotes the total number of unallocated XPUs within the rack (Silberschatz et al., 2018). In our simulator, we fully allocate the XPUs in a cluster by sampling sub-block slice sizes from the publicly available production distribution (Jouppi et al., 2023) and then deallocate 20% of the slices at random. Figure 3(d) shows fragmentation index across all 64 racks in a XPU cluster in our experiment. In the worst case, racks are susceptible to high fragmentation, as high as $I=0.7$, which hampers the ability to allocate slices to tenants, despite the availability of free accelerators.
3.3. L3: High blast radius of accelerator failures
@@ 5. Morphlux’s software orchestrator
  • fault manager: fault manager reacts to an accelerator failure by identifying an unallocated healthy chip within the same rack and logically replacing the failed chip, addressing the limitation in §3.3. This in-place replacement of faulty chips is possible by connecting the healthy chip via optical circuits to the remaining slice.
@@ 5.2. Fragmented slice allocator
Hardware topology. We denote the rack with graph $G^{\prime}:\langle S,I\rangle$. $S$ is the set of physical servers in a rack and $I$ are the optical fibers connecting consecutive servers. P(u,v){P}(u,v) is the set of pre-computed paths between physical servers $u$ and $v$, and can be sampled with a variety of algorithms (Yen, 1970).
@@ 5.3. In-place fault tolerance with Morphlux
Probability $\geq K$ failures. We define $Z(K)$ as the probability that $\geq K$ SRGs fail in $N$ SRGs. Let $p_{i}$ be the $P_{\text{fail}}$ of the $i^{th}$ SRG, then $\sum_{i=K}^{N}\sum_{\begin{subarray}{c}A\subseteq\{1,\ldots,N\}\\ |A|=i\end{subarray}}\left[\prod_{j\in A}p_{j}\right]\left[\prod_{j\notin A}(1-p_{j})\right]. Z(K)Z(K)$ is expensive to compute as it iterates over all subsets of sizes $K$ to $N$.
Key insight: Our key insight is that the inner sum of $Z(K)$ can be represented as a recursive relation with smaller sub-problems. Let $dp[i][k]$ be the probability of $k$ failures in $i$ SRGs, then the recursive relation is given by $dp[i][k]=dp[i-1][k-1]*(p_{i})+dp[i-1][k]*(1-p_{i})$. We formulate this as a dynamic program which computes larger dpsdps using smaller solutions as shown in Figure 5(a). Given the $dp$ matrix, we can compute Z(K) by summing over all probabilities where at least $K$ failures occur, i.e., $Z(K)=\sum_{k=K}^{N}dp[N][k]$. This reduces the computational complexity from $O(2^{N})$ to polynomial $O(N^{2})$, making it tractable for large networks with many SRGs. We pick $K$ SRGs such that $Z(K)\leq(1-S)$ where $S$ is $\frac{SLO}{100}$. This $K$ guarantees that the probability of $K$ or more SRGs failing is less than (1S)(1-S). So, all the failures that occur during the remaining $S$ fraction of the time can be gracefully handled by adding $K$ more SRGs.
@@ Appendix A Bandwidth Underutilization examples
The multi-dimensional bucket algorithm sequentially executes data transfers in a ring across all the dimensions of the torus. So, in a 3D torus, 3 rings are executed in the order of dimensions $XYZ$. As a result, connectivity in two of the three dimensions is always underutilized since only one ring is active at a given time. Intermediate computation prevents overlapping the three rings to prevent the interconnect under-utilization. To address this under-utilization, researchers have proposed algorithms that subdivide the data buffer to execute several multidimensional bucket algorithms in different sequences of dimensions, e.g., $YZX$ and $ZXY$, simultaneously such that all the dimensions are utilized throughout the collective (Sensi et al., 2024). However, this does not offer better performance, as we will discuss later.
@@
Impact of the underutilized bandwidth on $\alpha-\beta$ costs. We use the $\alpha-\beta$ cost model (Shah et al., 2023) to reason about the cost of collective communication. $\alpha$ cost is the software overhead of sending data buffers. $\beta$ cost is the transmission delay, which is inversely proportional to link bandwidth. The AllReduce bucket ring algorithm on a $D$ dimensional torus has $D$ ReduceScatter operations followed by $D$ AllGather operations and can reach optimal $\beta$ cost with simultaneous rings in all $D$ dimensions. However, Table 2 shows that the cost of one ReduceScatter on Slice-1 is proportional to $3$ times the optimal ($\frac{1}{bandwidth}$), because the slice is utilizing the bandwidth of only one dimension of the torus.
@@
$G^{\prime}:\langle L,T\rangle$: Slice request graph with $L$ slots and $T$ edges. $G:\langle S,I\rangle$: Physical rack’s graph with $S$ servers and $I$ edges. P(u,v){P}(u,v): Set of paths between servers $u$ and $v$, $\forall u,v\in{S}$, where each $Q\in{P}(u,v)$ is a list of edges $i\in{I}$

### v3

@@ 1. Introduction
$1.$ Overcoming Semiconductor Packaging Barriers
@@ 2. Torus-based datacenter fabrics for ML
Advantages of torus fabrics. In these datacenters, accelerators are directly connected to each other in a torus-shaped topology, avoiding packet switches. The resulting direct-connect torus-based fabrics offer two key advantages. First, direct or point-to-point connectivity between accelerators eliminates delays due to network contention if the fabric topology matches the communication pattern of the collective communication primitive. And second, well-known collective communication algorithms offer optimal performance on certain contention-free torus topologies (Wu et al., 2023; Jouppi et al., 2023; Zhao et al., 2024; Sack and Gropp, 2015a). For example, the torus topology works well with multi-dimensional bucket ring algorithms (Sack and Gropp, 2015a; Jouppi et al., 2023; Jain and Sabharwal, 2010; Sack and Gropp, 2015b). This algorithm sequentially executes data transfers in a ring across all the dimensions of the torus (XYZXYZ). Ring-based algorithms have an accelerator communicate with only two other accelerators at a given time, making communication in a ring on a direct-connect torus congestion-free.
@@ 3.1. L1: Bandwidth under-utilization
In a direct-connect torus topology, congestion occurs when multiple data transfers occur simultaneously on the same link (Leighton et al., 1994; Rothvoss, 2012). Accelerators in a slice can fully use their links in a dimension only when there is no congestion in that dimension. By design, an accelerator in a 3D torus has congestion-free access to links in all three dimensions only if the slice spans an entire rack or multiple racks. Consider Figure 3(a) with multiple sub-rack slices. Notice the links between slices 44 and 33 in the Z dimension. If both tenant slices use these links, they will cause congestion on them. To avoid causing congestion, these slices must avoid using them. As a result, the bandwidth of accelerators in each slice is underutilized by 33% because the slices have access to only 2 of the 3 dimensions. Smaller slices can have up to 66% lower bandwidth, as we show with examples in the Appendix A.
@@ 3.2. L2: Compute fragmentation
Specifically, allocations and deallocations of sub-block slices across the photonic datacenter fabric over time leads to fragmentation of resources within a rack. We quantify the fragmentation index II of a rack as: $I=1-\frac{S}{T}$, where $S$ is the number of XPUs in the largest allocatable slice, and $T$ denotes the total number of unallocated XPUs within the rack (Silberschatz et al., 2018). In our simulator, we fully allocate the XPUs in a cluster by sampling sub-block slice sizes from the publicly available production distribution (Jouppi et al., 2023) and then deallocate 20% of the slices at random. Figure 3(d) shows fragmentation index across all 64 racks in a XPU cluster in our experiment. In the worst case, racks are susceptible to high fragmentation, as high as $I=0.7$, which hampers the ability to allocate slices to tenants, despite the availability of free accelerators.
$3.$ L3: High blast radius of accelerator failures
@@ 5. Morphlux’s software orchestrator
  • fault manager: fault manager reacts to an accelerator failure by identifying an unallocated healthy chip within the same rack and logically replacing the failed chip, addressing the limitation in §$3.$ This in-place replacement of faulty chips is possible by connecting the healthy chip via optical circuits to the remaining slice.
@@ 5.2. Fragmented slice allocator
Hardware topology. We denote the rack with graph $G^{\prime}:\langle S,I\rangle$. $S$ is the set of physical servers in a rack and II are the optical fibers connecting consecutive servers. P(u,v){P}(u,v) is the set of pre-computed paths between physical servers $u$ and $v$, and can be sampled with a variety of algorithms (Yen, 1970).
@@ 5.3. In-place fault tolerance with Morphlux
Probability $\geq K$ failures. We define $Z(K)$ as the probability that $\geq K$ SRGs fail in NN SRGs. Let $p_{i}$ be the $P_{\text{fail}}$ of the $i^{th}$ SRG, then $\sum_{i=K}^{N}\sum_{\begin{subarray}{c}A\subseteq\{1,\ldots,N\}\\ |A|=i\end{subarray}}\left[\prod_{j\in A}p_{j}\right]\left[\prod_{j\notin A}(1-p_{j})\right]. Z(K)Z(K)$ is expensive to compute as it iterates over all subsets of sizes $K$ to NN.
Key insight: Our key insight is that the inner sum of $Z(K)$ can be represented as a recursive relation with smaller sub-problems. Let $dp[i][k]$ be the probability of $k$ failures in $i$ SRGs, then the recursive relation is given by $dp[i][k]=dp[i-1][k-1]*(p_{i})+dp[i-1][k]*(1-p_{i})$. We formulate this as a dynamic program which computes larger dpsdps using smaller solutions as shown in Figure 5(a). Given the dpdp matrix, we can compute Z(K) by summing over all probabilities where at least $K$ failures occur, i.e., $Z(K)=\sum_{k=K}^{N}dp[N][k]$. This reduces the computational complexity from $O(2^{N})$ to polynomial $O(N^{2})$, making it tractable for large networks with many SRGs. We pick $K$ SRGs such that $Z(K)\leq(1-S)$ where $S$ is $\frac{SLO}{100}$. This $K$ guarantees that the probability of $K$ or more SRGs failing is less than (1S)(1-S). So, all the failures that occur during the remaining $S$ fraction of the time can be gracefully handled by adding $K$ more SRGs.
@@ Appendix A Bandwidth Underutilization examples
The multi-dimensional bucket algorithm sequentially executes data transfers in a ring across all the dimensions of the torus. So, in a 3D torus, 3 rings are executed in the order of dimensions XYZXYZ. As a result, connectivity in two of the three dimensions is always underutilized since only one ring is active at a given time. Intermediate computation prevents overlapping the three rings to prevent the interconnect under-utilization. To address this under-utilization, researchers have proposed algorithms that subdivide the data buffer to execute several multidimensional bucket algorithms in different sequences of dimensions, e.g., YZXYZX and ZXYZXY, simultaneously such that all the dimensions are utilized throughout the collective (Sensi et al., 2024). However, this does not offer better performance, as we will discuss later.
@@
Impact of the underutilized bandwidth on $\alpha-\beta$ costs. We use the $\alpha-\beta$ cost model (Shah et al., 2023) to reason about the cost of collective communication. $\alpha$ cost is the software overhead of sending data buffers. $\beta$ cost is the transmission delay, which is inversely proportional to link bandwidth. The AllReduce bucket ring algorithm on a $D$ dimensional torus has $D$ ReduceScatter operations followed by $D$ AllGather operations and can reach optimal $\beta$ cost with simultaneous rings in all $D$ dimensions. However, Table 2 shows that the cost of one ReduceScatter on Slice-1 is proportional to 33 times the optimal ($\frac{1}{bandwidth}$), because the slice is utilizing the bandwidth of only one dimension of the torus.
@@
$G^{\prime}:\langle L,T\rangle$: Slice request graph with $L$ slots and $T$ edges. $G:\langle S,I\rangle$: Physical rack’s graph with $S$ servers and II edges. P(u,v){P}(u,v): Set of paths between servers $u$ and $v$, $\forall u,v\in{S}$, where each $Q\in{P}(u,v)$ is a list of edges $i\in{I}$